OpenStudy (anonymous):

For this problem, assume the box contains 9 red balls, 4 white balls, and 3 blue balls, and that we choose two balls at random from the box. What is the probability of neither being red given that neither is white?

OpenStudy (dumbcow):

Let A be case where both balls are not red Let B be case where both balls are not white A and B is the case where neither is red nor is white, both are blue P(A | B) = P(A and B)/P(B) P(A) = 7/16 * 6/15 = 7/40 P(B) = 12/16 * 11/15 = 11/20 P(A and B) = 3/16 *2/15 = 1/40 --> P(A |B) = 1/40 * 20/11 = 1/22